(1 point) Sketch S, the part of the plane - 10x + 5y + 10z = -39 selected by the inequality x2 + y2 + z2 <9. Assume a uniformSolve part b: Note the answer should be a vector
Answer #1

Electric field due to surface charge, E =  LON 250

ps = Surface charge density

E= 36 + 10-4 2* 8.854 * 10-12 = 2.03 * 103\widehat{a_{N}} N/C.

When the plane is z=2.6, then the normal direction will be az. That is \widehat{a_{N}} = \overrightarrow{a_{z}}

E= 2.03 * 103 \overrightarrow{a_{z}} N/C

Since origin is below z=2.6, the electric field at origin E(0) will be negative of the net electric field due to given surface charge.

E(0) =  - 2.03 * 103 \overrightarrow{a_{z}} N/C